Call for Papers: TPHOLs’99 The 12th International Conference on Theorem Proving in Higher Order Logics

The 1999 International Conference on Theorem Proving in Higher Order Logics will be the twelfth in a series that dates back to 1988. The Conference will be held on the four days Tuesday 14 September - Friday 17 September 1999 at l’Université de Nice, on the French Riviera. As with previous TPHOLs conferences, the 1999 formal proceedings will be published in Lecture Notes in Computer Science (LNCS). More information is available from the conference web site

    http://www-sop.inria.fr/croap/TPHOLs99/

Topic

The program committee welcomes submissions on all aspects of theorem proving in higher order logics and on applications. of higher order theorem proving. This includes, but is not limited to

• hardware verification and synthesis

• verification of security and communications protocols

• software refinement, transformation and verification

• industrial applications of theorem provers

• advances in theorem prover technology

• comparisons of various approaches to theorem proving

• proof automation and decision procedures

• incorporation of theorem provers into larger systems

• exploitation of external tools within theorem provers

• onnecting theorem provers together

• user interfaces for theorem provers

• development and extension of higher order logics

The scope of the conference encompasses all mechanised reasoning tools for higher order logics.

Submission and Publication

Submissions are invited in the following categories:

Formal Research Papers: Submissions in this category will be fully refereed, and those accepted will be published in a volume of the Springer-Verlag series Lecture Notes in Computer Science, which will be available at the conference. Authors of accepted papers are expected to present their work at the conference.

Work in Progress Papers: Submissions in this category will not be formally refereed, but their content and relevance will be reviewed. Accepted papers will be published in an INRIA technical report which will be available at the conference. Authors of accepted papers are expected to present a brief outline of their work at the conference, and to prepare a poster for display at the conference venue.

Submissions under both categories must be kept to 18 pages or less, and should be prepared according to the LNCS style. Papers may be submitted electronically in either PostScript or Portable Document Format (PDF). Complete instructions on how to submit are available from the conference web site. Please read these before submitting your paper.

Important Dates

Deadline for research paper submissions:  February, 28th 1999
Research paper acceptance notification:  April, 15th 1999
Camera-ready copy for research papers due:  May, 12th 1999
Deadline for progress paper submissions:  May, 19th 1999
Progress paper acceptance notification:  June, 30th 1999
Camera-ready copy for progress papers due:  July, 25th 1999
Conference:  September, 14-17 1999

Program Committee

• Mark Aagaard (Intel)

• Sten Agerholm (IFAD)

• David Basin (Freiburg)

• Yves Bertot (INRIA)

• Richard Boulton (Edinburgh)

• Gilles Dowek (INRIA)

• Mike Gordon (Cambridge)

• Jim Grundy (ANU)

• Elsa Gunter (Lucent)

• Joshua Guttman (Mitre)

• John Harrison (Intel)

• Doug Howe (Lucent)

• Bart Jacobs (Nijmegen)

• Sara Kalvala (Warwick)

• Tom Melham (Glasgow)

• Paul Miner (NASA)

• Malcolm Newey (ANU)

• Tobias Nipkow (Munich)

• Sam Owre (SRI)

• Christine Paulin-Mohring (LRI)

• Lawrence Paulson (Cambridge)

• Sofiène Tahar (Concordia)

Conference Organisation

The TPHOLs’99 conference organisers are Yves Bertot (INRIA), Gilles Dowek (INRIA), André Hirschowitz (Université de Nice), Christine Paulin-Mohring (LRI) and Laurent Théry (INRIA). They may be contacted at

                                                                               TPHOLs99@sophia.inria.fr

Versions of this Call for Papers are available from the conference web site in ASCII, DVI, HTML, PDF and PostScript.